The time taken by a particle performing SHM on a straight line to pass from a point A to B where its velocities are same is 2 seconds .After another 2 seconds it returns to B .What is the time period of oscillation in seconds??

Grade:11

1 Answers

Arun
5 years ago
According to the question, a and b points are such that they are located at same distances from the equilibrium position. Here also it is said that velocity is same ,i.e; not only the magnitude but also the directions are same. So, total time period of oscillation = 2×( time taken to go from a to b + the next time taken to return at b) = 2×(2+2)= 8 sec(ans)